Sol Summary – April 19th

Crew 280 Sol Summary Report 19-04-2023

Sol: 3

Summary title: Indoor Adventures for a Windy Day on Mars

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: The Hypatia I crew woke up with high spirits and a renewed sense of adventure. As the crew began their day, some members attended a yoga class taught by Ariadna Farrés (Health & Safety Officer), while others had the chance to take their first shower in days. The refreshing feeling of clean skin was a welcome change from the usual routine of sponge baths and hygiene wipes!

Once everyone was ready for the day, the crew gathered on the upper deck to have breakfast and attend a group meeting. Originally, the Hypatia I crew had planned to perform an EVA to Compass Rock with three main goals in mind: (i) collect sand samples for Laia Ribas’ (GreenHab Officer) CatMart project, (ii) evaluate the feasibility of different navigation strategies developed by Ariadna Farrés (Scientist & HSO); and (iii) test our own Midland radios for long-distance communication. However, as they checked the forecast, they realized that high winds were expected throughout the day. Therefore, the crew decided to be conservative and cancel their EVA. It was a tough call, as they were excited to work on their experiments outside the habitat. But safety always comes first on Mars, and they did not want to risk the well-being of anyone on the mission.

As the afternoon progressed, the crew took advantage of their time and focused on their experiments inside the habitat. At some point, the heater stopped working, causing the Hab to become quite cold. Cesca Cufí (Crew Engineer) was able to troubleshoot this problem with the assistance of Mission Support, but it is now failing again. We hope to be able to fix it, as temperatures will continue to drop during the night.

Despite the unexpected challenge, the crew remained positive, looking forward to a relaxing evening ahead. They will have homemade bread for dinner and are now excited to wind down with a delicious meal and a movie, ready for whatever challenges that tomorrow might bring.

Look ahead plan: We are planning on conducting two EVAs and continuing our experimental work.

Anomalies in work: Non-functional robotic telescope and Spirit needs a new battery wire.

Weather: Sunny, but high winds. Low 37F/High 53F.

Crew physical status: Nominal

EVA: Not applicable today.

Reports to be filed: Sol Summary, EVA request, HSO Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: Replacement of stove knobs, if possible.

Journalist Report – April 19th

Crew 280 Journalist Report 19-04-2023

Author: Núria Jar, Crew Journalist

The Hypatia crew members are getting into the swing of things on Mars. They have a structured daily schedule, divided into 30-minute slots. It is crucial in maintaining productivity and morale in such a challenging environment. Today was the first day all astronauts were on time, which suggests that they are adapting well to the new Martian routine.

The first group activity of the day is breakfast at 9AM. Prior to this, half of the crew is allowed to do some exercise and take a very quick shower. They have been successfully managing their water consumption, with an average of 11 liters per person, which equals 2,9 gallons. Interestingly, the plants in the base consume more water, at 48 liters per day, which is around 12,6 gallons.

The GreenHab Officer, Laia Ribas, has already planted the first seeds in order to get sprouts of bean mix and salad mix, which includes alfalfa, radish, broccoli and clover. The crew members have already eaten some delicious vegetables from the GreenHab, such as one cucumber, little tomatoes, and some lettuce and mint.

Fresh vegetables help provide a sense of normalcy in an otherwise isolated and challenging environment. The same applies for other Earthly pleasures, such as freshly baked bread. The Crew Scientist, Ariadna Farrés Basiana, spent around three hours baking bread for all crew members to lift the spirits of the team.

The third day on Mars was an indoors day, as there was no extravehicular activity (EVA) due to high winds. The crew focused on their experiments and brainstormed their outreach projects. Discussing science popularization strategies can be just as important as conducting experiments, especially in a simulated mission where sharing the results with the public can help inspire future scientists. It’s especially crucial to inspire young girls to pursue careers in STEM fields and break down gender barriers in science.

Overall, it was a productive and cold day despite the weather conditions. It is quite concerning that tomorrow’s temperature is expected to be around -3 Celsius in the morning, but regardless of the cold weather on Mars, Hypatia is choosing to stay warm indoors.

GreenHab Report – April 19th

Crew 280 GreenHab Report 20-04-2023
GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 68.4 F

Hours of supplemental light: N/A

Daily water usage for crops: 9.5 gallons

Daily water usage for research and/or other purposes: 1.5 gallons (Scientific Dome)

Water in Blue Tank: 158 gallons

Time(s) of watering for crops: 10.30, 17:45 h

Changes to crops: those dried leaves of tomato plants have been removed. Sticks have been added to two tomato plants. The tomato plant next to the door has been moved to another place to be protected from the wind because the plant looked not happy.

General watering of plants and crops was done today. There was wind and the temperature was not high, therefore the door was closed the whole day. Salad seeds for the crew have been cropped.

Harvest: 118 g yellow and 145 red tomatoes, respectively and 3 g of mint

Support/supplies needed

Operations Report – April 19th

Crew 280 Operations Report 19-04-2023

SOL: 3

Name of person filing report: Cesca Cufí (Crew Engineer) and Neus Sabaté (Mission Specialist)

Non-nominal systems:

Robotic observatory. Notes: replacement parts are scheduled to be installed during crew 261 rotation.

Rover Spirit. Notes: corroded battery wire. Replacement pending (Sergii). Rover can only be used for short distance EVA.


Spirit rover used: no

Hours: 216.7

Beginning charge: –

Ending charge: –

Currently charging: no

Opportunity rover used: no

Hours: 120.7

Beginning charge: –

Ending charge: –

Currently charging: no

Curiosity rover used: no

Hours: 229.8

Beginning charge: –

Ending charge: –

Currently charging: no

Perseverance rover used: no

Hours: 261.1

Beginning charge: –

Ending charge: –

Currently charging: no

General notes on rovers: none

Summary of Hab operations:

Water use: 20 gal (from 9AM Sol2 to 8AM Sol3)

Water (static tank): 443 gal (at 6:30PM). Water tank refilled this afternoon.

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: used from 18:00 to 21:00 for personal communication and reports.

Summary of suits and radios:

Suits: not used.

Radios: EVA radios not used.Station radios

All radio and spacesuit batteries checked before EVA: ok.

Summary of GreenHab operations:

Water use: 1,5 gal for experiments in the Science Dome

Heater: off

Supplemental light: off

Harvest: 3g of mint and 263g of tomatoes

Summary of Science Dome operations:

Laia Ribas (Green Hab Officer) continued preparing her experiments.

Neus Sabaté (Mission Specialist) started her experiment.

Dual split: off

Summary of RAM operations:

Cesca Cufí (Crew Engineer) continued her project integration.

Summary of any observatory issues: none

Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.

Summary of health and safety issues: we have replaced batteries of portable gas detectors. The exhausted batteries have been placed in a ziplog bag and left next to the B and NB bins in the lower deck.

Summary of other system’s issues:

Hab heater: it turned off about 2PM. No reason found. Crew Engineer and Commander troubleshoot the issue with the help of mission control. The heater showed the following Light code message: one short blink followed by 3 long blinks. The following actions were executed following mission control orders. The heater was turned off and on again, without results. Then the heater switch on the voltage panel was turned off and on again. The heater LED indicator was then constantly on, meaning that the heater was receiving 24VAC. After a few seconds, the heater turned on.

At 6.45PM the heater turned off again and it was troubleshot the same way.

Some bridles of the corridor tents were missing or broken after overnight winds. We replaced the damage with new bridles.

Questions, concerns and requests to Mission Support:

Spirit rover: replacement of then corroded wire

Kitchen: knobs replacement

Science dome: we have been organizing the science dome in order to fit all the planned experiments of the mission. Yesterday we had to move the microwave from one of the tables in order to have more dedicated space for biology experiments. Today, we allocated it in its original place but we realized that that space was needed to run the genome sequencing experiment. Therefore, we request to Mission Support the permission to store the microwave in a safe place during the mission and put it back to its original location once the experiments are completed. Attached you will find a pdf that explains how the different available locations of the science dome will be used and the suggested temporary location for the microwave.

Copyright © The Mars Society. All rights reserved. | Main Site